which is the paralel line?
(0,4),5y - 15x= 6
answers
y=-1/3x+3
y=3x+4
y=-1/3+4
y=3x-4
y=-1/4+3
which one passes through the point .select the correct answer.
y=4x+6
y=5x+2
y=5x+9
y=2x+5
y=6x+2


the graphs of pair of equations are parallel pertingular or neither
8x=8y=1
8x-8y=8

Answer by stanbon(48558)

which is the parallel line?
(0,4),5y - 15x= 6
Solve for y: y=3x+(6/5)
The slope is 3
y=3x+4 and y=3x-4 are parallel to the given line
-----------------------------------
which one passes through the point .select the correct answer.
Answer: none of those listed below.
y=4x+6
y=5x+2
y=5x+9
y=2x+5
y=6x+2
==================
the graphs of pair of equations are parallel perpendicular or neither
Answer: Parallel because they both have slope = 8/8=1
8x=8y=1
8x-8y=8
--------------
Cheers,
Stan H.